OpenStudy (anonymous):

2 times x over 6 times x plus 7 equal 2 over x. What is the value of x?

OpenStudy (amistre64):

you having to practice words to math notation or what?

OpenStudy (dumbcow):

great equation :) whats the order of operations? 2x/(6x+7) = 2/x ??

OpenStudy (anonymous):

no kinda like radical exspressions

OpenStudy (anonymous):

and yes dumbcow

OpenStudy (dumbcow):

i think you mean rational expressions

OpenStudy (anonymous):

what is the value of x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i really need help

OpenStudy (dumbcow):

well you have 2 ratios equal to each other...cross-multiply 2x^2 = 12x +14

OpenStudy (dumbcow):

divide by 2 and set equal to 0 x^2 -6x -7 = 0 Factor and solve

OpenStudy (anonymous):

soo the answer is?

OpenStudy (dumbcow):

... the roots of the quadratic above ^^ you have done these before

OpenStudy (anonymous):

not reallly im have to do this assignment and its due today

OpenStudy (anonymous):

i dnt understand tht well

OpenStudy (dumbcow):

i showed you the steps to get it into a quadratic. use the quadratic formula, do you learn how to use that?

OpenStudy (anonymous):

yes but wats the answer

OpenStudy (dumbcow):

you tell me, and i'll tell you if its right. quadratic formula: x = -b +-sqrt(b^2-4ac) / 2a